# Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations

### v3

@@ 7.1.4. OTIF
We compare the run-time of the video processor of Spatialyze with that of OTIF on object tracking. We run each of the two systems on the sampled nuScenes datasets, and compare the number of frames processed per second for each system. OTIF’s training is 61m37s. After it is trained, OTIF tracks object at 17.3 FPS. For Spatialyze, our video processor applies all optimizations elaborated in § 6, and it processes videos at 18.3-39.5 FPS to get the tracks for each query. Without counting OTIF’s training time, Spatialyze still tracks objects faster.
@@ 7.1.5. SkyQuery
SkyQuery is an aerial drone sensing platform for detecting and tracking vehicle from a top-down videos. A user inputs an aerial drone’s video and its geospatial data, then constructs a query to detect and track objects. Spatialyze and SkyQuery both process geospatial videos by 1) detecting objects of interest, 2) finding the objects’ 3D locations, and then 3) tracking the objects.
To compare with SkyQuery, we implemented the 3 steps mentioned as Spatialyze’s video processing steps, and used SkyQuery’s customized YOLOv3 for car detection, their 3D location estimator, and SORT (Bewley et al., 2016) for object tracking. We retained the rest of Spatialyze’s components as well as the optimization steps. Then, we executed both of the systems on Q10 on SkyQuery’s video and geospatial metadata, and compared execution times. As shown in Fig. 12, Spatialyze processes 6.08 FPS, while SkyQuery processes 5.15 FPS, making us 18% faster. We write 15 lines of S-Flow vs 4 lines of SkyQuery to express Q10. The majority of the lines in S-Flow are spent to setup the world and inputting videos and geospatial metadata. While we input videos and geospatial metadata into Spatialyze using S-Flow, we need to specify such input directories in the command line when starting SkyQuery.
As our evaluation uses the same ML models as SkyQuery, our speedup is entirely due to Spatialyze’s leveraging of the query’s semantics to prune video frames and avoid executing ML functions on them. Specifically, Spatialyze uses the Road Visibility Pruner to exclude video frames that do not contain any cycling lanes. We do not evaluate our accuracy in this evaluation as we use the exact same ML inferences functions as those in SkyQuery. As the only optimization technique applied in this evaluation is the Road Visibility Pruner, it only prunes out video frames that do not contain a cycling lane, and thus, does not affect the query result.
Figure 12. Video frames processed per second for each system.
@@ 7.2. Ablation Study
To evaluate each of our optimization techniques, we manually construct video processing plans for the following 7 experiment setups. (SB) Baseline is a Baseline plan without any optimization. It includes the 4 processing steps mentioned discussed in § 5.2.2. (S1) is the Baseline plan with the Road Visibility Pruner. (S2) is the Baseline plan with the Object Type Pruner. (S3) is the Baseline plan with Monocular Depth Estimator replaced with the 0pt. (S4) is the Baseline plan with the Exit Frame Sampler. (S5) enables all optimizations except the Exit Frame Sampler. (S6) enables all optimizations. We evaluated Spatialyze on Q1-4 because Spatialyze can apply all or most of the optimization techniques to the executions of such queries; therefore, we can compare the effects of each optimization technique.
@@ 7.2.1. Execution time
Figure 13. (Above) Average video processing time for each video of 20 seconds, comparing each optimization technique. (Below) AssA against (SB) vs. other experiment setups.
@@ 7.2.2. Accuracy
In Fig. 13, the Road Visibility Pruner (S1) causes a small AssA drop of 0.4% for Q3-4 and 4.7% for Q1-2. The Object Type Pruner (S2) also causes a small AssA drop of 2.5% from Q2-4 and 5.3% for Q1. We do not include (S3) into the results since the object tracker only concerns the 2D bounding boxes of objects, 0pt does not affect the AssA of the object tracker. With all 3 optimization techniques (S5), the video processor still produces accurate object tracks of 93.4% on average compared to (SB), while speeding up 2.5-5.3$\times$ of its runtime. With all optimization techniques (S6), the video processor’s accuracy for object tracking is 15.5% on average, while speeding up 1.65-4.32%, additionally.
@@ 8. Conclusion
We presented Spatialyze, a system for geospatial video analytics that leverages geospatial metadata and physical properties of objects in videos to speed up video processing. Spatialyze can process videos of arbitrary length by streaming each frame through the operators in workflows expressed using S-Flow. Spatialyze’s DSL allows users to declaratively specify their geospatial video workflows, and our evaluation shows Spatialyze’s efficiency in processing such workflows when compared to SOTA systems.

### v4

@@ 7.1.4. OTIF
We compare the run-time of the video processor of Spatialyze with that of OTIF on object tracking. We run each of the two systems on the sampled nuScenes datasets and compare the number of frames processed per second for each system. OTIF’s training is 61m37s. After it is trained, OTIF tracks objects at 17.3 FPS. For Spatialyze, our video processor applies all optimizations elaborated in § 6, and it processes videos at 18.3-39.5 FPS to get the tracks for each query. Without counting OTIF’s training time, Spatialyze still tracks objects faster.
@@ 7.1.5. SkyQuery
SkyQuery is an aerial drone sensing platform for detecting and tracking vehicles from top-down videos. A user inputs an aerial drone’s video and its geospatial data, then constructs a query to detect and track objects. Spatialyze and SkyQuery both process geospatial videos by 1) detecting objects of interest, 2) finding the objects’ 3D locations, and then 3) tracking the objects.
To compare with SkyQuery, we implemented the 3 steps mentioned as Spatialyze’s video processing steps, and used SkyQuery’s customized YOLOv3 for car detection, their 3D location estimator, and SORT (Bewley et al., 2016) for object tracking. We retained the rest of Spatialyze’s components as well as the optimization steps. Then, we executed both of the systems on Q10 on SkyQuery’s video and geospatial metadata, and compared execution times. As shown in Fig. 5, Spatialyze processes 6.08 FPS, while SkyQuery processes 5.15 FPS, making us 18% faster. We write 15 lines of S-Flow vs 4 lines of SkyQuery to express Q10. The majority of the lines in S-Flow are spent on setting up the world and inputting videos and geospatial metadata. While we input videos and geospatial metadata into Spatialyze using S-Flow, we need to specify such input directories in the command line when starting SkyQuery.
As our evaluation uses the same ML models as SkyQuery, our speedup is entirely due to Spatialyze’s leveraging of the query’s semantics to prune video frames and avoid executing ML functions on them. Specifically, Spatialyze uses the Road Visibility Pruner to exclude video frames that do not contain any cycling lanes. We do not evaluate our accuracy in this evaluation as we use the same ML inference functions as those in SkyQuery. As the only optimization technique applied in this evaluation is the Road Visibility Pruner, it only prunes out video frames that do not contain a cycling lane, and thus, does not affect the query result.
[…]
Runtime comparison of Spatialyze against OTIF, NuScenes Devkit, EVA, VIVA (both their data and NuScenes), and SkyQuery (with their data). Spatialyze is faster in all cases except when running Q8 compared to EVA
The bar chart above shows the runtime breakdown of each streaming operator, comparing between SB, S1, S2, S3, S4, S5, S6, and for Q1, Q2, Q3, Q4. The chart below shows the AssA against SB of S1, S2, S4, S5, S6, and for Q1, Q2, Q3, Q4.
[…]
@@ 7.2. Ablation Study
To evaluate each of our optimization techniques, we manually construct video processing plans for the following 7 experiment setups. (SB) Baseline is a Baseline plan without any optimization. It includes the 4 processing operators discussed in § 5.2.2. (S1) is the Baseline plan with the Road Visibility Pruner. (S2) is the Baseline plan with the Object Type Pruner. (S3) is the Baseline plan with the Monocular Depth Estimator replaced with the 0pt. (S4) is the Baseline plan with the Exit Frame Sampler. (S5) enables all optimizations except the Exit Frame Sampler. (S6) enables all optimizations. We evaluated Spatialyze using Q1-4 because Spatialyze can apply all or most of the optimization techniques for the execution of such queries; therefore, we can compare the effects of each optimization technique.
@@ 7.2.2. Accuracy
In Fig. 5, the Road Visibility Pruner (S1) causes a small AssA drop of 0.4% for Q3-4 and 4.7% for Q1-2. The Object Type Pruner (S2) also causes a small AssA drop of 2.5% from Q2-4 and 5.3% for Q1. We do not include (S3) into the results since the object tracker only concerns the 2D bounding boxes of objects; 0pt does not affect the AssA of the object tracker. With all 3 optimization techniques (S5), the video processor still produces accurate object tracks of 93.4% on average compared to (SB), while speeding up 2.5-5.3$\times$ of its runtime. With all optimization techniques (S6), the video processor’s accuracy for object tracking is 84.5% on average, while speeding up 1.65-4.32%, additionally.
@@ 8. Conclusion
We presented Spatialyze, a system for geospatial video analytics that leverages geospatial metadata and physical properties of objects in videos to speed up video processing. Spatialyze can process videos of arbitrary length by streaming each frame through the operators in workflows expressed using S-Flow. Spatialyze’s DSL allows users to specify their geospatial video workflows declaratively, and our evaluation shows Spatialyze’s efficiency in processing such workflows when compared to SOTA systems.
Acknowledgements.
We thank Daniel Kang, Francisco Romero, Favyen Bastani, Edward Kim, and our anonymous reviewers for their insightful feedback. This work is supported in part by NSF grants IIS-1955488, IIS-2027575, ARO W911NF2110339, ONR N00014-21-1-2724, and DOE awards DE-SC0016260 and DE-SC0021982.
{CJK*}
UTF8gbsn